An electron is a subatomic particle (m = 9.11 10-31 kg) that is subject to electric forces. An electron moving in the +x direction accelerates from an initial velocity of +5.46 105 m/s to a final velocity of +1.93 106 m/s while traveling a distance of 0.041 m. The electron's acceleration is due to two electric forces parallel to the x axis: vector F 1 = +8.15 10-17 N, and vector F 2, which points in the -x direction. Find the magnitudes of the net force acting on the electron and the electric force vector F 2.
Answer #1

here

use the work - energy theorem :

W = deltaKE

Fnet * S = 0.5 * m (vf^2 - vi^2)

Fnet * 0.041 = 0.5 * 9.11 * 10^-31 ((1.93 * 10^6 )^2 - ( 5.46 * 10^5)^2)

Fnet = 3.8 * 10^-17 N

then

the electric force ;

Fnet = 8.15 * 10-17 - F

3.8 * 10^-17 = 8.15 * 10-17 - F

F = 8.15 * 10-17 - 3.8 * 10^-17 = 4.35 * 10^-17 N